A cat chases a mouse across a 0.98 m high
table. The mouse steps out of the way, and
the cat slides off the table and strikes the floor
1.0 m from the edge of the table.
The acceleration of gravity is 9.81 m/s
2
.
What was the cat’s speed when it slid off
the table?
Answer in units of m/s
1
Expert's answer
2019-10-08T10:01:34-0400

The time of flight is


"t=\sqrt{\frac{2h}{g}}"

The cat’s speed when it slid off the table:


"v=\frac{x}{t}=x\sqrt{\frac{g}{2h}}"

"v=1.0\sqrt{\frac{9.81}{2(0.98)}}=2.2\frac{m}{s}"